How far in cm would a light travel through diamond in 1.00 the index of refraction of diamond is 2.419


1
Expert's answer
2021-01-13T11:35:27-0500
"s=vt=\frac{c}{n}t\\s=\frac{3\cdot10^8}{2.419}1=1.24\cdot10^8\ m"